a train travelling at 20km/hr is approaching a platform a bird is siiting on a pole on the platform when the train is at a distance of 2 km from the pole brakes are applied which produce a uniform deceleration in it at that instant the bird flies towards the train at 60km/hr and after touching the nearest point on the train flies back to the pole and then flies towards the train and continues repeating itself calculate how much distance will the bird have flown before the train stopps

Grade:11

4 Answers

Arun Kumar IIT Delhi
9 years ago
Hi
Since the train has to stop at the station
So



Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty

palak singhal
9 years ago
Alok mishra
5 years ago
The above answer is 100%wrong the correct answer is :
V2=u2+2as
0=400+2×a×2
Hence a=-100km/hr2
Now,v=u+at
T=1/5sec
Now the time bird would fly is equal to time travel by train so v×t=distance covered by bird
So, 1/5×60=distance traveled by bird
=12km
Mana Ranjan Mishra
5 years ago
The above mentioned answer is right but the time calulacalc above is written as 1/5sec. That should be written as 1/5 hr.